# Quantum Universality in Composite Systems: A Trichotomy of Clifford Resources

### v1

@@ III A Geometric Criterion for Infiniteness
Proving that a finitely generated matrix group is infinite is, in general, a non-trivial algebraic problem [detinko2013deciding]. While the irreducibility of the adjoint action ensures that the group is not trapped in a proper Lie subgroup of lower dimension, it does not rule out the possibility that the group is finite. To address this, we introduce a topological condition based on the discrete nature of finite groups.
@@
Since we are interested in subgroups of $\mathrm{SU}(d)$, but physical operations are equivalent up to global phases, the appropriate topological setting is the projective unitary group $\text{PU}(d)=\mathrm{SU}(d)/Z(\mathrm{SU}(d))\cong\mathrm{U}(d)/\mathrm{U}(1)$. To capture the proximity to the identity in this quotient space, we employ a projective operator distance.
@@
$\mathrm{dist}_{\mathrm{proj}}(U,I)\coloneqq\min_{\phi\in[0,2\pi)}\|e^{i\phi}U-I\|,$ (8)
@@
Note that the minimization is well-defined because the function $f(\phi)=\|e^{i\phi}U-I\|$ is continuous on the compact interval $0,2\pi]$. Also, while $U\in\mathrm{SU}(d)$, the optimal representative $e^{i\phi}U$ may lie in $\mathrm{U}(d)$.
@@ III.1 Proof of the Criterion
We now rely on a structural property established in the proof of Theorem 5.7 in Ref. Dixon1971: for any finite subgroup $G^{\prime}\subset\mathrm{GL}(d,\mathbb{C})$, the subgroup $A=\langle\{g\in G^{\prime}:\|g-I\|<1/2\}\rangle$ is a normal abelian subgroup. Since $h^{\prime}\in G^{\prime}$ satisfies $\|h^{\prime}-I\|<1/2$, it implies $h^{\prime}\in A$, so $A$ is non-trivial. However, since the adjoint representation of $G^{\prime}$ (which coincides with that of $G$) is irreducible, Lemma 11 forces any normal abelian subgroup to be central. Thus, $h^{\prime}$ must be a scalar matrix. This implies $h$ is a scalar, which contradicts the assumption that $\mathrm{dist}_{\mathrm{proj}}(h,I)>0$. Therefore, $G$ cannot be finite. ∎
@@ III.2 Spectral Characterization via Projective Distance
To evaluate the projective distance efficiently, we first introduce a geometric measure of the angular concentration of the eigenvalues.
@@ IV Case I: Prime Dimensions
We begin our analysis with the simplest case: dimensions $d=p$ where $p$ is a prime number. In this setting, the Clifford group exhibits a remarkable structural rigidity. To understand this, we must first establish a general correspondence between the group’s action on the Lie algebra and the symplectic geometry of the discrete phase space. This framework will prove crucial not only for establishing universality in prime dimensions but also for diagnosing the obstructions that arise in prime-power dimensions (Sec. V).
@@ IV.1 The Adjoint Action and $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ Orbits
To explicitly decompose the adjoint representation of the Clifford group $C(d)$ on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ into a direct sum of irreducible subrepresentations, we employ the theory of representations of group extensions developed by Clifford [clifford1937, Serre1977].
@@ IV.2 Irreducibility of the Adjoint Action
The first condition of Theorem 4 is the irreducibility of the action on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$. We can now prove this characterization using the tools developed.
@@
Conversely, if $d$ is composite, there exists a proper divisor $k$ ($1<k<d$). The set of vectors with $\gcd(u,d)=k$ forms an orbit disjoint from the set of vectors with $\gcd(u,d)=1$. This implies the existence of at least a proper non zero invariant subspaces, rendering the representation reducible. ∎
@@ IV.3 Maximality and Universality
To establish universality for prime dimensions $d=p$, we rely on the fact that any extension of the Clifford group by a non-Clifford gate forces the group to become infinite. This property stems from the maximality of the Clifford group within the finite subgroups of the projective unitary group. While this fact is widely recognized in quantum information [Gottesman1999], a rigorous mathematical proof relies on deep results from finite group theory.
For the general case of systems with $n$-qudits ($d$ prime), Nebe, Rains, and Sloane [Nebe2001] established maximality using sophisticated number-theoretic techniques. Specifically, they utilized $l$-adic integrality and reduction modulo suitable rational primes to apply the classification of subgroups of finite general linear groups. However, as noted in Ref. [Nebe2001], for the single-qudit case (d=pd=p)—which is central to this work—the sufficiency of the primality condition follows from earlier work by Lindsey [Lindsey1970]. Unlike the general multi-qudit case, Lindsey’s proof relies primarily on character theory and the classification of linear groups of prime degree. Complementing this classical result, we demonstrate that primality is also a necessary condition, thereby establishing the following full characterization. For completeness, we provide a self-contained proof of the forward direction based on a Corollary in [Lindsey1970].
@@
For the remaining small primes $p\in\{2,3,5\}$, maximality follows from the explicit classification of finite primitive linear groups of degree $p$. If a finite extension $G\supsetneq\mathcal{C}_{p}$ existed, it would necessarily possess a Sylow $p$-subgroup of order at least $|\mathcal{C}_{p}|_{p}\geq p^{4}$ (accounting for the Heisenberg group and $SL(2,\mathbb{Z}/p\mathbb{Z})$ part). However, the classifications by Blichfeldt [Blichfeldt1917] and Brauer [Brauer1967] reveal that no primitive group—other than the normalizers of extraspecial groups—supports such a structure. For instance, for $p=5$, Brauer established that the only other primitive candidates involve central extensions of $A_{5},A_{6},$ or $\text{PSL}(2,11)$, whose Sylow $5$-subgroups have orders at most $25$, strictly less than the required order. Similar order constraints on the Sylow $p$-subgroups rule out extensions for $p=2$ and $p=3$.
@@
IV.4 Universality and Diagonal Resources
@@
This corollary shifts the burden of finding universal sets to simply identifying a single non-Clifford gate. The most natural candidates are diagonal gates.
Definition 21.
Definition 22.
$\delta_{\zeta}(x,y)=\frac{\zeta(x+y)}{\zeta(x)\zeta(y)}.$ (14)
Lemma 23.
@@
The conjugation action on the shift operator $X^{y}$ is given by:
This lemma provides a constructive method to generate universal gate sets. For a prime dimension $p$, any diagonal gate whose differential is not a bicharacter (corresponding to a non-quadratic phase) is universal.
IV.5 The $T_{s}$-gate Family
The criteria for universality derived in the previous sections—maximality for prime dimensions and irreducibility via mixing for composite dimensions—share a common requirement: the availability of a unitary that lies strictly outside the local Clifford group. To characterize such operators, for any dimension $d$, we introduce a generic family of diagonal gates.
Definition 24 (The $T_{s}$-gate Family).
This family generalizes the standard Clifford hierarchy. For instance, $T_{d}$ recovers the Pauli $Z$ gate (order $d$), and $T_{2d}$ corresponds to the generalized phase gate.
To determine the “Cliffordness” of $T_{s}$ for an arbitrary dimension, we analyze its conjugation action on the Pauli shift operator $X$. A direct calculation yields:
For $T_{s}$ to belong to the Clifford group $\mathcal{C}_{d}$, it must normalize the Heisenberg group. Since the result is proportional to the operator $X$, this condition holds if and only if the $e^{2\pi i/s}I$ is in the Heisenberg group:
  • Odd Dimensions: The phases are generated by the primitive $d$-th root of unity $\omega=e^{2\pi i/d}$. Thus, $T_{s}\in\mathcal{C}_{d}\iff s\mid d$.
  • Even Dimensions: The Clifford group generates phases of order $2d$. Specifically for qubits ($d=2$), the group includes the order-4 phase $i$. Thus, generally, $T_{s}\in\mathcal{C}_{d}\iff s\mid 2d$.
The previous discussion leads to a unified criterion for when $T_{s}$ constitutes a non-Clifford resource.
Proposition 25.
Corollary 26.
Let $p$ be a prime number. The set $\mathcal{C}_{p}\cup\{T_{s}\}$ is single-qudit universal if and only if $s\nmid K_{p}$.
V Case II: Prime Power Dimensions
While prime dimensions provide a rigid structure where the Clifford group is maximal and irreducible, prime power dimensions $d=p^{m}$ (with $m\geq 2$) introduce a rich internal structure derived from the arithmetic of the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$. In this regime, the two primary pillars of universality no longer hold: the adjoint action becomes reducible, and the group loses its maximality.
@@ V.1 Reducibility and Orbit Decomposition
In Sec. IV, we established that the irreducibility of the Clifford group’s action is equivalent to the transitivity of the $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ group on the Pauli indices (Lemma 15). While the prime case ($d=p$) guarantees transitivity, the prime power case ($d=p^{m}$, $m\geq 2$) introduces structural modifications due to the existence of non-trivial ideals in the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$.
@@
For a prime power dimension $d=p^{m}$ ($m\geq 2$), the adjoint representation of the Clifford group on $\mathfrak{sl}(d,\mathbb{C})$ is reducible. The algebra decomposes into a direct sum of $m$ irreducible orthogonal invariant subspaces:
@@
This result fundamentally shifts the criteria for universality compared to the prime case. Since the Lie algebra is fragmented into irreducible subspaces $\mathcal{W}_{k}$ that the Clifford group cannot connect, a supplementary gate must do more than simply generate infinite order; it must actively “mix” these subspaces, bridging irreducibility.
@@ V.2 Orbit-mixing Gates and Irreducibility
$\operatorname{Ad}_{T_{\zeta}}(X^{a})=\zeta(a)X^{a}D_{a},$ (20)
where $D_{a}$ is the diagonal operator with entries $(D_{a})_{xx}=\delta_{\zeta}(a,x)$. Here, $\delta_{\zeta}$ is the differential previously introduced in Eq. (14).
@@
$\operatorname{Ad}_{T_{\zeta}}(X^{a})=\frac{\zeta(a)}{\sqrt{d}}\sum_{b\in\mathbb{Z}_{d}}\widehat{(\delta_{\zeta,a})}(b)V_{(a,b)}.$ (22)
@@
  • Case 1: $\mathcal{S}$ contains a subspace $\mathcal{W}_{k}$ with $k\geq 1$. Let $X^{p^{k}}\in\mathcal{W}_{k}\subseteq\mathcal{S}$. Since $\mathcal{S}$ is $G$-invariant, $\operatorname{Ad}_{T_{\zeta}}(X^{p^{k}})\in\mathcal{S}$. By Eq. (22) and the Orbit-Mixing Condition (Def. 28), the expansion of this vector contains a term $\beta V_{(p^{k},b)}$ with $\beta\neq 0$ and $b\in\mathbb{Z}_{d}^{\times}$. Since $b$ is a unit, $\gcd(p^{k},b,p^{m})=1$, which implies $V_{(p^{k},b)}\in\mathcal{W}_{0}$. Since $\mathcal{S}$ is invariant under the Clifford group (which acts transitively on the orbit associated with $\mathcal{W}_{0}$), containing one vector in $\mathcal{W}_{0}$ implies $\mathcal{W}_{0}\subseteq\mathcal{S}$.
@@ V.3 Universality via the $T_{s}$-gate Family
We now apply our general machinery—irreducibility via orbit-mixing gates and infiniteness via the geometric criterion—to the generic family of diagonal gates $T_{s}$. Recall that $T_{s}$ introduces phases proportional to the roots of unity of order $s$, see Eq. (15).

### v2

@@ III A Criterion for Infiniteness
Proving that a finitely generated matrix group is infinite is, in general, a non-trivial algebraic problem [13]. While the irreducibility of the adjoint action ensures that the group is not trapped in a proper Lie subgroup of lower dimension, it does not rule out the possibility that the group is finite. To address this, we introduce a topological condition based on the discrete nature of finite groups.
@@
Since we are interested in subgroups of $\mathrm{SU}(d)$, but physical operations are equivalent up to global phases, the appropriate topological setting is the projective unitary group $\mathrm{PU}(d)=\mathrm{SU}(d)/Z(\mathrm{SU}(d))\cong\mathrm{U}(d)/\mathrm{U}(1)$. To capture the proximity to the identity in this quotient space, we employ a projective operator distance.
@@
$\mathrm{dist}_{\mathrm{proj}}(U,I)\coloneqq\min_{\phi\in[0,1)}\|e^{2\pi i\phi}U-I\|,$ (14)
@@
Note that the minimization is well-defined because the function $f(\phi)=\|e^{2\pi i\phi}U-I\|$ is continuous on the compact interval $[0,1]$. Also, while $U\in\mathrm{SU}(d)$, the optimal representative $e^{2\pi i\phi}U$ may lie in $\mathrm{U}(d)$.
@@ III.1 Proof of the Criterion
We now rely on a structural property established in the proof of Theorem 5.7 in Ref. 14: for any finite subgroup $G^{\prime}\subset\mathrm{GL}(d,\mathbb{C})$, the subgroup $A=\langle\{g\in G^{\prime}:\|g-I\|<1/2\}\rangle$ is a normal abelian subgroup. Since $h^{\prime}\in G^{\prime}$ satisfies $\|h^{\prime}-I\|<1/2$, it implies $h^{\prime}\in A$, so $A$ is non-trivial. However, since the adjoint representation of $G^{\prime}$ (which coincides with that of $G$) is irreducible, Lemma 18 forces any normal abelian subgroup to be central. Thus, $h^{\prime}$ must be a scalar matrix. This implies $h$ is a scalar, which contradicts the assumption that $\mathrm{dist}_{\mathrm{proj}}(h,I)>0$. Therefore, $G$ cannot be finite. ∎
@@ III.2 Spectral Characterization via Projective Distance
To evaluate the projective distance efficiently, we first introduce a measure of the angular concentration of the eigenvalues.
@@ IV Case I: Prime Dimensions
We begin our analysis with the simplest case: dimensions $d=p$ where $p$ is a prime number. In this setting, the Clifford group exhibits a strong structural rigidity. To understand this, we must first establish a general correspondence between the group’s action on the Lie algebra and the symplectic geometry of the discrete phase space. This framework will prove useful not only for establishing universality in prime dimensions but also for diagnosing the obstructions that arise in prime-power dimensions (Sec. V).
@@ IV.1 The Adjoint Action and $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ Orbits
To explicitly decompose the adjoint representation of the Clifford group $C(d)$ on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ into a direct sum of irreducible subrepresentations, we employ the theory of representations of group extensions developed by Clifford [11, 36].
@@ IV.2 Irreducibility of the Adjoint Action
The first condition of the density criterion (Theorem 4) is the irreducibility of the action on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$. As formalized by Graydon et al[21], a finite group forms a unitary 2-design if and only if its adjoint representation on the space of operators decomposes into exactly two irreducible components: the identity and the traceless subspace.
While Graydon et al. demonstrated that this condition fails for composite dimensions (implying the Clifford group is not a 2-design in those cases), the following theorem establishes the positive counterpart: for prime dimensions, the adjoint action is strictly irreducible.
@@
Conversely, if $d$ is composite, there exists a proper divisor $k$ ($1<k<d$). The set of vectors with $\gcd(u,d)=k$ forms an orbit disjoint from the set of vectors with $\gcd(u,d)=1$. This implies the existence of at least one proper nonzero invariant subspace, rendering the representation reducible. ∎
@@ IV.3 Maximality and Universality
To establish universality for prime dimensions $d=p$, we rely on the fact that any extension of the Clifford group by a non-Clifford gate forces the group to become infinite. This property stems from the maximality of the Clifford group within the finite subgroups of the projective unitary group. While this fact is well-known in quantum information [20], a rigorous mathematical proof relies on deep results from finite group theory.
For the general case of systems with $n$-qudits ($d$ prime), Nebe, Rains, and Sloane [34] established maximality using number-theoretic techniques. Specifically, they utilized $l$-adic integrality and reduction modulo suitable rational primes to apply the classification of subgroups of finite general linear groups. However, as noted in Ref. [34], for the single-qudit case (d=pd=p)—which is central to this work—the sufficiency of the primality condition follows from earlier work by Lindsey [26]. Unlike the general multi-qudit case, Lindsey’s proof relies primarily on character theory and the classification of linear groups of prime degree. Complementing this classical result, we demonstrate that primality is also a necessary condition, thereby establishing the following full characterization. For completeness, we provide a self-contained proof of the forward direction based on a Corollary in [26].
@@
For the remaining small primes $p\in\{2,3,5\}$, maximality follows from the explicit classification of finite primitive linear groups of degree $p$. If a finite extension $G\supsetneq\mathcal{C}_{p}$ existed, it would necessarily possess a Sylow $p$-subgroup of order at least $|\mathcal{C}_{p}|_{p}\geq p^{4}$ (accounting for the Heisenberg group and $SL(2,\mathbb{Z}/p\mathbb{Z})$ part). However, the classifications by Blichfeldt [3] and Brauer [5] reveal that no primitive group—other than the normalizers of extraspecial groups—supports such a structure. For instance, for $p=5$, Brauer established that the only other primitive candidates involve central extensions of $A_{5},A_{6},$ or $\text{PSL}(2,11)$, whose Sylow $5$-subgroups have orders at most $25$, strictly less than the required order. Similar order constraints on the Sylow $p$-subgroups rule out extensions for $p=2$ and $p=3$.
@@
IV.4 Universality Resources in Prime Dimensions
@@
This corollary shifts the burden of finding universal sets to simply identifying a single non-Clifford gate. We can now identify specific resources based on the classifications established in the Preliminaries.
First, regarding diagonal resources:
Corollary 28.
Let $p$ be a prime number. The set $\mathcal{C}_{p}\cup\{T_{s}\}$ is single-qudit universal if and only if $s\nmid K_{p}$ (as defined in Proposition 15).
Second, we consider permutation resources. Since the Clifford group contains only affine permutations (Proposition 9), any permutation that violates the affine condition $x\mapsto ax+b$ constitutes a universal resource.
Corollary 29.
Let $p\geq 5$ be a prime. Any non-affine permutation $\pi\in S_{p}$ yields a universal gate set $\mathcal{C}_{p}\cup\{P_{\pi}\}$. In particular, the simple transposition $\tau=(0\;1)$, which swaps the first two basis states and fixes the rest, is universal for all $p\geq 5$.
@@
By Corollary 27, it suffices to show $P_{\tau}\notin\mathcal{C}_{p}$, which holds if $\tau$ is not affine. Suppose $\tau(x)=ax+b$. Evaluated at $x=0$, $\tau(0)=1\implies b=1$. Evaluated at $x=1$, $\tau(1)=0\implies a(1)+1=0\implies a=-1$. Evaluated at $x=2$ (since $p\geq 3$, the element 2 exists), $\tau(2)=2$. However, the affine formula predicts $\tau(2)=(-1)(2)+1=-1$. Thus, equality holds only if $2\equiv-1\pmod{p}$, i.e., $3\equiv 0\pmod{p}$. This is possible only for $p=3$. For all primes $p\geq 5$, the transposition is non-affine and therefore universal. ∎
V Case II: Prime-power Dimensions
While prime dimensions provide a rigid structure where the Clifford group is projectively finite maximal and irreducible, prime-power dimensions $d=p^{m}$ (with $m\geq 2$) introduce a structure derived from the arithmetic of the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$. In this case, the two primary pillars of universality no longer hold: the adjoint action becomes reducible, and the group loses its maximality.
To achieve universality, one must add a gate that restores irreducibility while ensuring infiniteness. We show that two classes of non-Clifford gates accomplish this: diagonal gates from the $T_{s}$ family (Sec. V.3) and non-affine permutations such as transpositions (Sec. V.4).
@@ V.1 Reducibility and Orbit Decomposition
In Sec. IV, we established that the irreducibility of the Clifford group’s action is equivalent to the transitivity of the $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ group on the Pauli indices (Lemma 22). While the prime case ($d=p$) guarantees transitivity, the prime-power case ($d=p^{m}$, $m\geq 2$) introduces structural modifications due to the existence of non-trivial ideals in the ring $\mathbb{Z}/{p^{m}}\mathbb{Z}$.
@@
For a prime-power dimension $d=p^{m}$ ($m\geq 2$), the adjoint representation of the Clifford group on $\mathfrak{sl}(d,\mathbb{C})$ is reducible. The algebra decomposes into a direct sum of $m$ irreducible orthogonal invariant subspaces:
@@
This result changes the criterion for universality compared to the prime case. Since the Lie algebra is fragmented into irreducible subspaces $\mathcal{W}_{k}$ that the Clifford group cannot connect, a supplementary gate must do more than simply generate infinite order; it must actively “mix” these subspaces, bridging irreducibility.
@@ V.2 Orbit-mixing Gates and Irreducibility
$\operatorname{Ad}_{T_{\zeta}}(X^{a})=X^{a}D_{a},$ (21)
where $D_{a}$ is the diagonal operator with entries
$(D_{a})_{xx}=\frac{\zeta(a+x)}{\zeta(x)}.$ (22)
Equivalently, $(D_{a})_{xx}=\zeta(a)\,\delta_{\zeta}(a,x)$, where $\delta_{\zeta}$ is the 1-coboundary previously introduced in Eq. (10).
@@
$\operatorname{Ad}_{T_{\zeta}}(X^{a})=\frac{1}{\sqrt{d}}\sum_{b\in\mathbb{Z}_{d}}\widehat{(\delta_{\zeta,a})}(b)V_{(a,b)}.$ (24)
@@
  • Case 1: $\mathcal{S}$ contains a subspace $\mathcal{W}_{k}$ with $k\geq 1$. Let $X^{p^{k}}\in\mathcal{W}_{k}\subseteq\mathcal{S}$. Since $\mathcal{S}$ is $G$-invariant, $\operatorname{Ad}_{T_{\zeta}}(X^{p^{k}})\in\mathcal{S}$. By Eq. (24) and the Orbit-Mixing Condition (Def. 31), the expansion of this vector contains a term $\beta V_{(p^{k},b)}$ with $\beta\neq 0$ and $b\in\mathbb{Z}_{d}^{\times}$. Since $b$ is a unit, $\gcd(p^{k},b,p^{m})=1$, which implies $V_{(p^{k},b)}\in\mathcal{W}_{0}$. Since $\mathcal{S}$ is invariant under the Clifford group (which acts transitively on the orbit associated with $\mathcal{W}_{0}$), containing one vector in $\mathcal{W}_{0}$ implies $\mathcal{W}_{0}\subseteq\mathcal{S}$.
@@ V.3 Universality via the $T_{s}$-gate Family
We now apply our general machinery—irreducibility via orbit-mixing gates and infiniteness via the infiniteness criterion—to the generic family of diagonal gates $T_{s}$. Recall that $T_{s}$ introduces phases proportional to the roots of unity of order $s$, see Eq. (11).
Note that $T_{s}$ is a unitary diagonal operator (hence $T_{s}\in\mathrm{U}(d)$), and it may fail to have determinant one. Since our density and infiniteness criteria are projective (and the adjoint action is insensitive to global phases), we may freely replace $T_{s}$ by any scalar multiple in $\mathrm{SU}(d)$ without affecting the arguments.
@@
V.4 Permutations as Non-Clifford Resources
While diagonal gates of the $T_{s}$ family provide a method to achieve universality, permutation gates offer an alternative class of non-Clifford resources. In this subsection, we analyze when non-affine permutations can supplement the Clifford group to generate a dense subgroup of $\mathrm{SU}(d)$.
V.4.1 Induced Diagonal Operators
Conjugation by a permutation matrix transforms the Pauli-$Z$ operator into a diagonal operator that encodes the permutation structure. Explicitly, the adjoint action yields the diagonal operator $D_{\pi}$ given by:
$P_{\pi}ZP_{\pi}^{\dagger}=D_{\pi},\quad\text{where}\quad(D_{\pi})_{yy}=\omega^{\pi^{-1}(y)}.$ (30)
Since $D_{\pi}$ and $Z$ are related by unitary conjugation, they share the same spectrum: the complete set of $d$-th roots of unity. Consequently, $D_{\pi}$ has maximal spectral span $\ell(D_{\pi})=(d-1)/d$ and cannot directly satisfy the infiniteness criterion of Theorem 17.
To extract a useful resource, we consider the difference operator that isolates the deviation from the identity permutation.
Definition 34.
For a permutation $\pi\in S_{d}$, the difference operator is:
$h_{\pi}=D_{\pi}Z^{\dagger}=P_{\pi}ZP_{\pi}^{\dagger}Z^{\dagger},$ (31)
with diagonal entries $(h_{\pi})_{yy}=\omega^{\pi^{-1}(y)-y}$.
The difference operator measures how much $\pi$ displaces each index. For an affine permutation $\pi(x)=ax+b$, we have $\pi^{-1}(y)=a^{-1}(y-b)$, so the exponent becomes:
$\pi^{-1}(y)-y=(a^{-1}-1)y-a^{-1}b,$
which is linear in $y$. Thus, $h_{\pi}\in H(d)$ for affine permutations. For non-affine permutations, $h_{\pi}$ exhibits a non-linear phase structure.
V.4.2 Transpositions and the Infiniteness Criterion
We analyze the simplest non-affine permutations: transpositions.
Proposition 35.
Let $\tau=(j\;k)$ be the transposition swapping indices $j$ and $k$ with $j<k$. The difference operator is:
$h_{\tau}=\mathrm{diag}(1,\ldots,1,\omega^{k-j},1,\ldots,1,\omega^{-(k-j)},1,\ldots,1),$ (32)
where the entries $\omega^{k-j}$ and $\omega^{-(k-j)}$ occur at positions $j$ and $k$, respectively. The spectral span is:
$\ell(h_{\tau})=\frac{2(k-j)}{d}.$ (33)
Proof.
Since transpositions are self-inverse, $\tau^{-1}=\tau$. The diagonal entries of $h_{\tau}$ are $(h_{\tau})_{yy}=\omega^{\tau(y)-y}$:
  • For $y=j$: $\tau(j)=k$, so $(h_{\tau})_{jj}=\omega^{k-j}$.
  • For $y=k$: $\tau(k)=j$, so $(h_{\tau})_{kk}=\omega^{j-k}=\omega^{-(k-j)}$.
  • For $y\notin\{j,k\}$: $\tau(y)=y$, so $(h_{\tau})_{yy}=1$.
The eigenvalues are $\{1,\omega^{k-j},\omega^{-(k-j)}\}$ with multiplicities d2d-2, 11, and $1$, respectively. These span an arc from angle $2\pi(k-j)/d$ to $2\pi(k-j)/d$, giving spectral span $\ell(h_{\tau})=2(k-j)/d$. ∎
Corollary 36.
The transposition $\tau=(j\;k)$ generates an infinite group $\langle\mathcal{C}_{d},P_{\tau}\rangle$ provided:
$k-j<\frac{d\cdot\arcsin(1/4)}{\pi}\approx 0.0805\cdot d.$ (34)
In particular, the adjacent transposition $\tau=(0\;1)$ satisfies this bound for all $d\geq 13$.
Proof.
By Lemma 20, the projective distance is:
$\mathrm{dist}_{\mathrm{proj}}(h_{\tau},I)=2\sin\left(\frac{\pi\ell(h_{\tau})}{2}\right)=2\sin\left(\frac{\pi(k-j)}{d}\right).$
The infiniteness criterion (Theorem 17) requires this distance to be less than $1/2$:
$2\sin\left(\frac{\pi(k-j)}{d}\right)<\frac{1}{2}\implies\sin\left(\frac{\pi(k-j)}{d}\right)<\frac{1}{4}.$
For the adjacent transposition with kj=1k-j=1, this becomes $\sin(\pi/d)<1/4$, which holds when $d>\pi/\arcsin(1/4)\approx 12.43$. ∎
V.4.3 Irreducibility Analysis for Transpositions
We now verify that the difference operator $h_{\tau}$ induced by the adjacent transposition satisfies the orbit-mixing condition of Definition 31.
Let $d=p^{m}$ with $m\geq 2$, and let $\tau=(0\;1)$. The phase function $\zeta:\mathbb{Z}_{d}\to\mathrm{U}(1)$ of the difference operator $h_{\tau}$ is given by $\zeta(0)=\omega$, $\zeta(1)=\omega^{-1}$, and $\zeta(y)=1$ for $y\geq 2$. For any proper divisor $u=p^{k}$ with $1\leq k\leq m-1$, a direct calculation of the differential row $f_{u}(y)=\zeta(u+y)/\zeta(y)$ yields:
$f_{u}(y)=\begin{cases}\omega^{-1}&y\in\{0,\,d-u+1\},\\ \omega&y\in\{1,\,d-u\},\\ 1&\text{otherwise},\end{cases}$ (35)
where indices are taken modulo $d$. Using this explicit form, we determine the spectral properties of the gate.
Proposition 37.
For $n\not\equiv 0\pmod{d}$, the Discrete Fourier Transform of the differential row $f_{u}$ is:
$\sqrt{d}\,\hat{f}_{u}(n)$ $=(\omega^{-1}-1)(1+\omega^{n(u-1)})$
$\quad+(\omega-1)(\omega^{-n}+\omega^{nu}).$ (36)
Proof.
The Fourier transform is $\sqrt{d}\,\hat{f}_{u}(n)=\sum_{y=0}^{d-1}f_{u}(y)\,\omega^{-ny}$. We decompose this sum by separating the deviation from the constant function:
$\sqrt{d}\,\hat{f}_{u}(n)=\sum_{y=0}^{d-1}\omega^{-ny}+\sum_{y\in S}(f_{u}(y)-1)\omega^{-ny},$
where $S=\{0,1,d-u,d-u+1\}$ is the support of the deviation.
For $n\not\equiv 0\pmod{d}$, the first sum vanishes: $\sum_{y=0}^{d-1}\omega^{-ny}=0$.
The second sum evaluates to:
$(\omega^{-1}-1)\cdot\omega^{0}+(\omega-1)\cdot\omega^{-n}$
$+(\omega-1)\cdot\omega^{-n(d-u)}+(\omega^{-1}-1)\cdot\omega^{-n(d-u+1)}.$
Using $\omega^{d}=1$, we simplify:
$\omega^{-n(d-u)}$ $=\omega^{-nd+nu}=\omega^{nu},$
$\omega^{-n(d-u+1)}$ $=\omega^{-nd+nu-n}=\omega^{n(u-1)}.$
Substituting and grouping terms yields Eq. (37). ∎
Proposition 38.
Let $d=p^{m}$ with $m\geq 2$. The difference operator $h_{\tau}$ for $\tau=(0\;1)$ is an orbit-mixing gate in the sense of Definition 31.
Proof.
We must show that for each proper divisor $u=p^{k}$ ($1\leq k\leq m-1$), there exists at least one unit $n\in\mathbb{Z}_{d}^{\times}$ such that $\hat{f}_{u}(n)\neq 0$. We prove that $n=1$ always works.
Suppose, for contradiction, that $\hat{f}_{u}(1)=0$. From Eq. (37) with $n=1$:
$(\omega^{-1}-1)(1+\omega^{u-1})+(\omega-1)(\omega^{-1}+\omega^{u})=0.$
Using the identity $\omega^{-1}-1=-\omega^{-1}(\omega-1)$ and dividing by $(\omega-1)\neq 0$:
$\omega^{-1}(1+\omega^{u-1})+(\omega^{-1}+\omega^{u})=0.$
Expanding and simplifying:
$\omega^{-1}-\omega^{u-2}+\omega^{-1}+\omega^{u}=0,$
which reduces to $\omega^{u}=\omega^{u-2}$, i.e., $\omega^{2}=1$.
Since $\omega=e^{2\pi i/d}$, this requires $d\mid 2$. For any prime power $d=p^{m}\geq 4$ with $m\geq 2$, this condition fails. Therefore, $\hat{f}_{u}(1)\neq 0$ for every proper divisor $u$, confirming the orbit-mixing condition. ∎
V.4.4 Universality Theorem for Transpositions
Combining the results on infiniteness and irreducibility, we obtain:
Theorem 39.
Let $d=p^{m}$ with $m\geq 2$. The group $G=\langle\mathcal{C}_{d},P_{\tau}\rangle$, where $\tau=(0\;1)$ is the adjacent transposition, is dense in $\mathrm{SU}(d)$.
Proof.
By Theorem 4, it suffices to verify irreducibility of the adjoint representation and infiniteness $G$.
Irreducibility: The difference operator $h_{\tau}=P_{\tau}ZP_{\tau}^{\dagger}Z^{\dagger}$ belongs to $G$. By Proposition 38, $h_{\tau}$ is an orbit-mixing gate. Theorem 32 then implies that $G$ acts irreducibly on $\mathfrak{sl}(d,\mathbb{C})$.
Infiniteness: For $d\geq 13$, Corollary 36 establishes that the projective distance satisfies $\mathrm{dist}_{\mathrm{proj}}(h_{\tau},I)<1/2$, and Theorem 17 implies that $G$ is infinite.
For the remaining prime-power cases $d\in\{4,8,9\}$, infiniteness has been verified by explicit computation using the computer algebra system Magma [4]. The following code confirms that $\langle\mathcal{C}_{d},P_{\tau}\rangle$ is infinite in each case:
1for d in [4, 8, 9] do
2 ord := LCM(d, 2*d);
3 F := CyclotomicField(ord);
4 z := F.1; omega := z^(ord div d);
5 zeta := z^(ord div (2*d));
6 H := (1/F!Isqrt(d)) *
7 Matrix(F,d,d,[omega^(j*k): j,k in [0..d-1]]);
8 P := DiagonalMatrix(F,
9 [zeta^(j*(j + d mod 2)): j in [0..d-1]]);
10 X := Matrix(F,d,d,[<i,(i mod d)+1,1>: i in [1..d]]);
11 tau := IdentityMatrix(F, d);
12 tau[1,1]:=0; tau[1,2]:=1; tau[2,1]:=1; tau[2,2]:=0;
13 G := sub<GL(d, F) | H, P, X, tau>;
[…]
15end for;
16
17// Output: d=4: Infinite=true
18// d=8: Infinite=true
19// d=9: Infinite=true
The two conditions of Theorem 4 are satisfied for all prime powers $d=p^{m}$ with $m\geq 2$, completing the proof. ∎